# EC-LDA : Label Distribution Inference Attack against Federated Graph Learning with Embedding Compression

### v1

@@ Abstract
Graph Neural Networks (GNNs) have been widely used for graph analysis. Federated Graph Learning (FGL) is an emerging learning framework to collaboratively train graph data from various clients. However, since clients are required to upload model parameters to the server in each round, this provides the server with an opportunity to infer each client’s data privacy. In this paper, we focus on label distribution attacks (LDAs) that aim to infer the label distributions of the clients’ local data. We take the first step to attack client’s label distributions in FGL. Firstly, we observe that the effectiveness of LDA is closely related to the variance of node embeddings in GNNs. Next, we analyze the relation between them and we propose a new attack named EC-LDA, which significantly improves the attack effectiveness by compressing node embeddings. Thirdly, extensive experiments on node classification and link prediction tasks across six widely used graph datasets show that EC-LDA outperforms the SOTA LDAs. For example, EC-LDA attains optimal values under both Cos-sim and JS-div evaluation metrics in the CoraFull and LastFM datasets. Finally, we explore the robustness of EC-LDA under differential privacy protection.
1 Introduction
1.1 Background
@@
Graph Neural Networks (GNNs), designed to process graph-structured data, have gained significant attention for their effectiveness across various applications including recommendation systems He et al. (2020), social networks Fan et al. (2019), and protein interaction prediction Jha et al. (2022). GNNs can capture information between neighboring nodes, enhancing the expressiveness of node embeddings and making them highly suitable for real-world applications.
@@
The performance of GNNs requires a large amount of data. However, due to privacy concerns and regulatory restrictions, machine learning platforms cannot access raw data directly, which makes centralized GNN learning challenging. In recent years, a lot of works He et al. (2021); Zhang et al. (2021) have integrated Federated Learning (FL) McMahan et al. (2017) with GNNs, proposing the Federated Graph Learning (FGL). FL is a distributed, privacy-preserving machine learning paradigm that enables clients to train models collaboratively while keeping their local data isolated. It addresses the challenge of data silo, where data is distributed across different sources and cannot be easily combined for joint analysis, by enabling model training without the need to share the raw data.
@@
However, recent studies have shown that local data in FL remains vulnerable to privacy attacks Zhu et al. (2019); Geiping et al. (2020); Zhao et al. (2020). One such attack is label distribution inference attacks (LDAs), which aim to infer the label distribution of a client’s local training data by analyzing the gradients shared between clients and the server. This represents a significant privacy threat in FL. For example, if multiple online shopping companies collaborate to train a recommendation system model (as shown in Figure 1), a malicious server with access to the label distribution of a private social network could target specific users, increasing the success rate of fraudulent activities and posing a serious threat to user security.
Figure 1: An overview of EC-LDA.
1.2 Previous Works and Problems
@@
Despite existing works on LDAs, there are significant challenges when applying them to FGL. Firstly, these attacks do have certain limitations. For example, Gu and Bai (2023); Wainakh et al. (2021) requires the use of real auxiliary datasets to execute the attacks. Yin et al. (2021) requires the use of non-negative activation functions. Additionally, the performance of attacks in Ma et al. (2023); Wainakh et al. (2021) declines as the local epochs in FL increase, achieving effective results only with a single local training epoch. These assumptions restrict their applicability in GNN scenarios. Furthermore, the message-passing characteristic among adjacent nodes in GNNs complicates node embeddings, potentially introducing irrelevant details into the model output, thereby making LDA in GNNs more challenging.
1.3 Our Contributions
@@
In this paper, we propose Embedding Compression-Label Distribution Inference Attack (EC-LDA), a novel LDA in FGL. Through exploring the factors behind the poor performance of LDAs in GNNs, we find there are strong correlation between the effectiveness of LDA and the variance of node embeddings. In GNNs, the variance in node embeddings primarily originates from the GNN layers and increases with the number of layers, which in turn degrades the attack performance of LDA. EC-LDA addresses this issue by compressing node embeddings, thereby reducing their variance and enhancing attack performance. Moreover, EC-LDA overcomes the drawbacks of existing methods. It keeps stable performance as training samples and local epochs increase and doesn’t need specific activation functions or extra datasets for attacks. Our main contributions are as follows:
@@
  • We apply EC-LDA to six graph datasets and conduct extensive experiments targeting both node classification and link prediction tasks. These experiments demonstrate that EC-LDA consistently achieves significant attack performance across various scenarios. EC-LDA achieves a Cos-sim as high as 1.000 under almost all cases.
  • We utilize DP-GNN (node-level DP) and label-DP (label-level DP) for local GNN training and evaluate the robustness of EC-LDA under these privacy protection mechanisms.
2 Preliminaries
2.1 Graph Neural Networks
In general, GNNs are designed to process data $G(V,\mathbb{E})$ structured as a graph. Here, $V$ represents the set of nodes and $\mathbb{E}$ represents the adjacency matrix of $G$. Each node $v_{i}\in V$ has a feature vector $u_{i}$. GNNs generate valuable node embeddings via message-passing, making them suitable for various downstream tasks like node classification and link prediction. In this paper, we consider Graph Convolutional Networks (GCN) Kipf and Welling (2016), Graph Attention Networks (GAT) Veličković et al. (2017), and GraphSAGE Hamilton et al. (2017) as target GNNs.
@@ 2.1 Graph Neural Networks
[…]
Variance*1000 of $I$ 41.052 23.719 21.283 10.640 4.547 2.031 0.817 0.224 0.067 0.003
$Err$ 4.367 3.433 2.765 2.217 1.523 1.012 0.599 0.309 0.138 0.028
Table 1: $Err$ with the variance of $I$ when local epochs $E$ is set to 1, the WikiCS dataset and a 2-layer GCN model are used.
2.2 Federated Graph Learning
A typical FGL system follows the FedAvg McMahan et al. (2017) algorithm. Specifically, the server sends an initial global model to all the clients. Then each client trains a local model with its local private data $G_{i}(V_{i},\mathbb{E}_{i})$ and shares its local model parameters with the server. The server then aggregates the local model parameters of all the clients to construct the global model’s parameters, which can be formulated as:
@@ 2.2 Federated Graph Learning
2.3 Threat Model
As shown in Figure 1, we consider the FGL scenario in which the server is malicious. In this scenario, the server is not only interested in accessing the label distributions of the clients’ private data, but also has the ability to manipulate the parameters of the deployed model. Furthermore, the server can analyze the gradients uploaded by the clients.
3 Related Work
The label distribution inference attack is a type of label recovery attack. Zhu et al. (2019) was the first to restore the training sample from gradients. They restored the input data and associated label from the gradients using gradient-matching. This method continuously optimizes the dummy input data and associated label by minimizing the mean square error of the gradients of the dummy sample with respect to the true gradients. Zhao et al. (2020) introduced iDLG and were the first to propose that with a non-negative activation function, privacy label can be extracted with 100% success rate from the signs of gradients at the output layer. Both of these methods are applicable only to single-sample training scenarios. Afterwards, Yin et al. (2021); Geng et al. (2021) extended the attacks to the mini-batch scenario, enhancing the applicability of the attacks.
Overall, most existing label recovery attacks focus on image datasets, and none of them discuss label distribution on graph datasets. In this paper, We focus on label distribution inference attacks against GNNs. More releated work please refer Appendix A.
4 Massage-Passing Issue in LDA
In this section, we will analyze how the message-passing mechanism in GNNs exacerbates LDA. We begin with an introduction to LDA. The prior analysis Geng et al. (2021) has shown that, for the $k$-th sample in a batch: $\Delta W_{l}^{(FC)}/E\approx\frac{1}{K}\sum_{k}(p_{k,l}\sum_{m}I_{k,m}-y_{k,l}\sum_{m}I_{k,m})$, where $\Delta W_{l}^{(FC)}$ is the sum of the gradients of the $l$-th output unit in the last fully connected layer along the input dimension, $I_{k,m}$ is the input of the $k$-th sample at the $m$-th input unit of the last fully connected layer, $p_{k,l}$ is the post-softmax probability at index $l$ of the $k$-th sample, $E$ is the local epochs. Furthermore, we have: $\sum_{k}y_{k,l}\cdot\bar{I}\approx\sum_{k}p_{k,l}I_{k}-K\Delta W_{l}^{(FC)}/E$, where $I_{k}=\sum_{m}I_{k,m}$, and $\bar{I}$ is the mean value of $I_{k}$.
@@ 4 Massage-Passing Issue in LDA
$d_{l}:=\sum_{k}y_{k,l}\approx\frac{\sum_{k}p_{k,l}I_{k}}{\bar{I}}-\frac{K\Delta W_{l}^{(FC)}}{E\bar{I}},$ (4)
@@
Figure 2: Illustrates the trend of variance of $I$ with the number of GCN layers. The experimental results at this point are the normalized results.
Figure 3: Illustration of the data flow during attack rounds in EC-LDA.
@@
$Err:=\frac{1}{K}\sum_{l}\lvert\sum_{k}y_{k,l}(I_{k}-\bar{I})\rvert=\frac{1}{K}\sum_{k}\lvert I_{k}-\bar{I}\rvert.$ (5)
We can intuitively see that if the variance of $I$ is smaller, the $Err$ is smaller and the performance of the attack is better, and vice versa. To verify our conjecture, we conduct experiments on FGL by performing attacks during mid-training round and analyze the variance of $I$ versus $Err$. Table 1 illustrates how $Err$ varies with the variance of $I$, demonstrating that the $Err$ declines as $I$’s variance declines, which confirms our suspicions.
As mentioned earlier, GNNs incorporate neighbor information through the message-passing mechanism. However, the message-passing mechanism also increases the variance of node embeddings. Moreover, the variance of the node embeddings increases with the number of GNN layers in the global model, and the node embeddings are directly related to $I$. We conduct experiments with GNN models featuring 1, 2, and 3 GCN layers and observe how the variance of $I$ changes with the number of GCN layers. Figure 2 illustrates the variance of $I$ as a function of the number of GCN layers. It can be seen that as the number of GCN layers increases, the variance of $I$ becomes larger. Therefore, minimizing the variance of $I$ is crucial for enhancing attack efficacy.
@@
Input: Number of clients $N$, global rounds $R$, attack rounds $A$
@@
5:        Generate dummy training data $DTD$
6:        Input $DTD$ into $W_{r-1}^{{}^{\prime}}$ to obtain $p_{k,l}$, $I_{k}$ and $\bar{I}$
@@
5 Our Methodology: EC-LDA
To minimize the variance of $I$, one effective approach is to constrain the absolute values of $I_{k}$ within a small range, which can be achieved by clipping the model parameters.
@@ 5 Our Methodology: EC-LDA
Based on the above analysis, we now describe the design of EC-LDA, an enhanced attack for GNNs. Unlike LDA, EC-LDA makes full use of Equation 4 by clipping the parameters of the global model before distributing it to clients, which reduces the variance of node embeddings, thus making the variance of $I$ smaller. As shown in Figure 3, EC-LDA consists of three main steps:
@@
  • Step 1: Clipping model. The server clips the initial model and sends it to clients to get the gradients. Algorithm 1 demonstrates the clipping method. First the $\ell_{2}$ norm $\mathbb{N}$ of the model is computed, and then $p$/max$(1,\frac{\mathbb{N}}{C})$, i.e., the clipped parameters, will replace the original parameters $p$.
  • Step 2: Forward propagation of dummy data. The server generates dummy training data randomly and inputs it into the initial model to obtain $I_{k}$, $\bar{I}$, and $p_{k,l}$.
  • Step 3: Calculating distribution. After the server gets gradients, $p_{k,l}$, $I_{k}$, and $\bar{I}$, Equation 4 is used to infer the label distributions of clients’ private data.
@@
Algorithm 2 demonstrates EC-LDA in FGL. The boxed part represents the additional component of EC-LDA compared to the normal training process. First, the server initializes the model parameters and attack results. If the server does not perform an attack in the current round $r$, the training process is no different from the normal training process. If the server is going to perform an attack in round $r$, it will save the initial model $W_{r-1}$ for round $r$ as $W_{r-1}^{{}^{\prime}}$ and clip $W_{r-1}$ with Algorithm 1, i.e. Step 1. The server generates dummy training data and feeds it into $W_{r-1}^{{}^{\prime}}$ to obtain $p_{k,l}$, $I_{k}$ and $\bar{I}$, i.e. Step 2. Then the server sends the clipped model down to all clients, and each client trains on its local private data and uploads the trained model. After training, server calculates the label distribution $D$ with Equation 4, and saves $D$ into $\mathbb{L}$, i.e. Step 3. The server uses the saved initial model $W_{r-1}^{{}^{\prime}}$ of round $r$ to replace $W_{r}$ so that the model trained and aggregated on the clipped model can be avoided to be passed to the next round. Therefore, a few attacks do not significantly impact model performance. The training process of clients does not differ from the general training process, clients first receive the model $W_{r-1}$ from the server, then train the model $W_{r-1}$ with the local privacy training data for $E$ times and upload the trained model to the server. In other words, EC-LDA trades one round of training resources for one round of label distributions of clients’ private training data.

### v3

@@ Abstract
Graph Neural Networks (GNNs) have been widely used for graph analysis. Federated Graph Learning (FGL) is an emerging learning framework to collaboratively train graph data from various clients. Although FGL allows client data to remain localized, a malicious server can still steal client private data information through uploaded gradient. In this paper, we for the first time propose label distribution attacks (LDAs [^1] The term “LDA” here is different from other machine learning terms like Latent Dirichlet Allocation.) on FGL that aim to infer the label distributions of the client-side data. Firstly, we observe that the effectiveness of LDA is closely related to the variance of node embeddings in GNNs. Next, we analyze the relation between them and propose a new attack named EC-LDA, which significantly improves the attack effectiveness by compressing node embeddings. Then, extensive experiments on node classification and link prediction tasks across six widely used graph datasets show that EC-LDA outperforms the SOTA LDAs. Specifically, EC-LDA can achieve the Cos-sim as high as 1.0 under almost all cases. Finally, we explore the robustness of EC-LDA under differential privacy protection and discuss the potential effective defense methods to EC-LDA. Our code is available at https://github.com/cheng-t/EC-LDA.
I Introduction
I-A Background
@@
Graph Neural Networks (GNNs), designed to process graph-structured data, have gained significant attention for their effectiveness across various applications including recommendation systems [1], social networks [2], and protein interaction prediction [3]. GNNs can capture information between neighboring nodes, enhancing the expressiveness of node embeddings and making them highly suitable for real-world applications.
@@
The performance of GNNs requires a large amount of data. However, due to privacy concerns and regulatory restrictions, machine learning platforms cannot access raw data directly, which makes centralized GNN learning challenging. In recent years, a lot of works [4, 5] have integrated Federated Learning (FL) [6] with GNNs, proposing the Federated Graph Learning (FGL). FL is a distributed, privacy-preserving machine learning paradigm that enables clients to train models collaboratively while keeping their local data isolated. It addresses the challenge of data silo, where data is distributed across different sources and cannot be easily combined for joint analysis, by enabling model training without the need to share the raw data.
@@
However, recent studies have shown that local data in FL remains vulnerable to various attacks [7, 8, 9, 10]. One such attack is label distribution inference attacks (LDAs) [11, 12], which aim to infer the label distribution of client-side local training data by analyzing the gradients shared between clients and the server. This represents a significant privacy threat in FL. For example, if multiple online shopping companies collaborate to train a recommendation system model (as shown in Fig. 1), a malicious server with access to the label distribution of a private social network could target specific users, increasing the success rate of fraudulent activities and posing a serious threat to user privacy.
Figure 1: Threat model of EC-LDA.
I-B Previous Works and Limitations
@@
Despite existing works on LDAs [11, 12, 13, 14], they all under non-GNN settings and have many limitations. For example, their attack methods require additional auxiliary datasets [11, 12] or specific activation functions [13] as assumption. Additionally, the performance of attack method declines as the local epochs in FL increase, achieving effective results only with a single local training epoch [14, 12]. These assumptions restrict their applicability to GNN scenarios. Furthermore, the message-passing characteristic among adjacent nodes in GNNs complicates node embeddings, potentially introducing irrelevant details into the model output, thereby making LDA in GNNs more challenging.
I-C Our Contributions
@@
In this paper, we propose Embedding Compression-Label Distribution Inference Attack (EC-LDA), a novel LDA in FGL. Through exploring the factors behind the poor performance of LDAs in GNNs, we find there are strong correlation between the effectiveness of LDA and the variance of node embeddings. In GNNs, the variance in node embeddings primarily originates from the GNN layers and increases with the number of layers, which in turn degrades the attack performance of LDA. EC-LDA addresses this issue by compressing node embeddings, thereby reducing their variance and enhancing attack performance. Moreover, EC-LDA overcomes the drawbacks of existing LDAs. It keeps stable performance as training samples and local epochs increase and doesn’t need specific activation functions or extra datasets for attacker. Our main contributions are as follows:
@@
  • We apply EC-LDA to six graph datasets and conduct extensive experiments targeting both node classification and link prediction tasks. These experiments demonstrate that EC-LDA consistently achieves significant attack performance across various scenarios. EC-LDA achieves a Cos-sim as high as 1.00 under almost all cases.
  • We utilize DP-GNN (node-level differential privacy) and Label-DP (label-level differential privacy) for local GNN training and evaluate the robustness of EC-LDA. We further discuss potential effective defense methods for EC-LDA.
I-D Paper Structure
The rest of the paper is organized as follows. Section II introduces the preliminaries knowledge and the threat model. In Section III, we analyze the message-passing in LDA and give the key observation. Though the findings, we proposed our attack EC-LDA approach in Section IV. The experimental results are presented in Section V. Section VI reviews related work, followed the limitations in Section VII. Lastly, conclusion in Section VIII.
II Preliminaries
II-A Graph Neural Networks
In general, GNNs are designed to process data $G(V,\mathbb{E})$ structured as a graph. Here, $V$ represents the set of nodes and $\mathbb{E}$ represents the adjacency matrix of $G$. Each node $v_{i}\in V$ has a feature vector $u_{i}$. GNNs generate valuable node embeddings via message-passing, making them suitable for various downstream tasks like node classification and link prediction. In this paper, we consider Graph Convolutional Networks (GCN) [15], Graph Attention Networks (GAT) [16], and GraphSAGE [17] as target GNNs.
@@ II-A Graph Neural Networks
where $u_{i}^{h-1}$ is the representation obtained at the (h1)(h\!-\!1)-th layer of node $v_{i}$, and $u_{i}^{0}$ is the node feature $u_{i}$ of node $v_{i}$, ${\mathbb{B}}_{i}$ represents the neighbors of node $v_{i}$, $AGG(\cdot)$ represents the aggregation function, $\sigma$ represents the activation function such as $ReLU$.
TABLE I: ErrErr with the variance of II when local epochs $E$ is set to 1 (WikiCS dataset and a 2-layer GCN model).
Variance*1000 of II 41.052 23.719 21.283 10.640 4.547 2.031 0.817 0.224 0.067 0.003
ErrErr 4.367 3.433 2.765 2.217 1.523 1.012 0.599 0.309 0.138 0.028
II-B Federated Graph Learning
A typical FGL system follows the FedAvg [6] algorithm. Specifically, the server sends an initial global model to all the clients. Then each client trains a local model with its local private data $G_{i}(V_{i},\mathbb{E}_{i})$ and shares its local model parameters with the server. The server then aggregates the local model parameters of all the clients to construct the global model’s parameters, which can be formulated as:
@@ II-B Federated Graph Learning
II-C Threat Model
As shown in Fig. 1, we consider the FGL scenario where a malicious server is curious about local private data and attempt to reconstruct the label distributions of the clients private data. For instance, consider multiple hospitals collaboratively training a GNN model to predict patients’ disease probabilities. Each hospital maintains local patient data graphs where nodes represent patients, edges correspond to medical relationships, and labels denote diagnosis results. By inferring hospital-specific label distributions, attackers could obtain sensitive regional health statistics, potentially revealing disease prevalence in specific healthcare jurisdictions.
Attacker’s Knowledge and Capability. In this scenario, the malicious server adversary can not only observe all gradient updates uploaded by clients but also actively tamper with global model parameters or the aggregation process [18, 19, 20].
III Message-Passing Issue in LDA
In this section, we will analyze how the message-passing mechanism in GNNs exacerbates LDA. We begin with an introduction to LDA. The prior analysis [21] has shown that, for the $k$-th sample in a batch: $\Delta W_{l}^{(FC)}/E\approx\frac{1}{K}\sum_{k}(p_{k,l}\sum_{m}I_{k,m}-y_{k,l}\sum_{m}I_{k,m})$, where $\Delta W_{l}^{(FC)}$ is the sum of the gradients of the $l$-th output unit in the last fully connected layer along the input dimension, $I_{k,m}$ is the input of the $k$-th sample at the $m$-th input unit of the last fully connected layer, $p_{k,l}$ is the post-softmax probability at index $l$ of the $k$-th sample, $E$ is the local epochs. Furthermore, we have: $\sum_{k}y_{k,l}\cdot\bar{I}\approx\sum_{k}p_{k,l}I_{k}-K\Delta W_{l}^{(FC)}/E$, where $I_{k}=\sum_{m}I_{k,m}$, and $\bar{I}$ is the mean value of $I_{k}$.
@@ III Message-Passing Issue in LDA
$d_{l}$ $=\sum_{k}y_{k,l}$ (4)
$\approx\frac{\sum_{k}p_{k,l}I_{k}}{\bar{I}}-\frac{K\Delta W_{l}^{(FC)}}{E\bar{I}},$
@@
Figure 2: Illustrates the trend of variance of II with the number of GCN layers. The experimental results at this point are the normalized results.
Figure 3: Illustration of the EC-LDA.
@@
$Err$ $=\frac{1}{K}\sum_{l}\lvert\sum_{k}y_{k,l}(I_{k}-\bar{I})\rvert$ (5)
$=\frac{1}{K}\sum_{k}\lvert I_{k}-\bar{I}\rvert$.
We can intuitively see that if the variance of $I$ is smaller, the $Err$ is smaller and the performance of the attack is better, and vice versa. To verify our conjecture, we conduct experiments on FGL by performing attacks during mid-training round and analyze the variance of $I$ versus $Err$. Table I illustrates how $Err$ varies with the variance of $I$, demonstrating that the $Err$ declines as $I$’s variance declines, which confirms our suspicions.
As mentioned earlier, GNNs incorporate neighbor information through the message-passing mechanism. However, the message-passing mechanism also increases the variance of node embeddings. Moreover, the variance of the node embeddings increases with the number of GNN layers in the global model, and the node embeddings are directly related to $I$. We conduct experiments with GNN models featuring 1, 2, and 3 GCN layers and observe how the variance of $I$ changes with the number of GCN layers. Fig. 2 illustrates the variance of $I$ as a function of the number of GCN layers. We scaled the variance to [0,1] through min-max normalization:
$V_{new}=\frac{V-V_{min}}{V_{max}-V_{min}},$ (6)
which $V$ is the variance of $I$. This normalization does not change the trend that the variance in all datasets increases with the number of GNN layers. The empirical results indicate a positive correlation between the depth of GCN layers and the variance of $I$. Consequently, optimizing attack efficacy necessitates minimizing the variance of $I$.
@@
Input: Client number $N$, global rounds $R$, attack rounds AA
@@
5:   Generate dummy training data $D_{dummy}$
6:   Input $D_{dummy}$ into $W_{r-1}^{{}^{\prime}}$ to obtain $p_{k,l}$, $I_{k}$ and $\bar{I}$
@@
IV Our Methodology: EC-LDA
Based on the above analysis, we conclude that the key to reducing error lies in minimizing the variance of $I$, which can be effectively achieved by constraining the absolute values of $I_{k}$ by clipping the model parameters.
IV-1 Overview
@@
We now provide the introduction to EC-LDA, an enhanced attack method targeting GNNs. Unlike LDA, EC-LDA makes full use of Equation 4 by clipping the parameters of the global model before distributing it to clients, which reduces the variance of node embeddings, thus making the variance of $I$ smaller.
As shown in Fig. 3, EC-LDA consists of three main steps:
@@
  • Step 1: Clipping model. The server clips the initial model and sends it to clients to get the gradients. Algorithm 1 demonstrates the clipping method. First the $\ell_{2}$ norm $\mathbb{N}$ of the model is computed, and then $p/max(1,\frac{\mathbb{N}}{C})$, i.e., the clipped parameters, will replace the original parameters $p$. After receiving the global model sent by the server, the client trains the model using its own private training data and uploads the trained model to the server.
  • Step 2: Forward propagation of dummy data. The server first generates dummy training data, which is then fed into the initial model. Through forward propagation, the server obtains intermediate outputs $_{k}, I¯\bar{I}$ from the input of the final fully-connected layer. Simultaneously, the model’s output is processed by the softmax layer to derive $p_{k,l}$.
  • Step 3: Calculating distribution. Following local client training and forward propagation of dummy training data, the server obtains intermediate variables gradients, $p_{k,l}$, $I_{k}$, and $\bar{I}$. It then calculates the private label distribution of client-side data by applying Equation 4.
IV-2 EC-LDA
@@
Algorithm 2 presents the server-side implementation of EC-LDA within the FGL framework. The boxed part represents the additional component of EC-LDA compared to the normal training process. First, the server initializes the model parameters and attack results (Line 1). If the server does not perform an attack in current round $r$, the training process is no different from the normal training process. Otherwise, the server saves the initial model $W_{r-1}$ for round $r$ as $W_{r-1}^{{}^{\prime}}$ and clip $W_{r-1}$ with Algorithm 1. The server generates dummy training data and feeds it into $W_{r-1}^{{}^{\prime}}$ to obtain $p_{k,l}$, $I_{k}$ and $\bar{I}$ (Line 4-6). Then the server sends the clipped model down to all clients, and each client trains on its local private data and uploads the trained model. After training, server calculates the label distribution $D$ with Equation 4, and saves $D$ into $\mathbb{L}$ (Line 11-12). The server uses the saved initial model $W_{r-1}^{{}^{\prime}}$ of round $r$ to replace $W_{r}$ (Line 17). This prevents propagation of clipping-affected models to subsequent rounds, thereby ensuring minimal performance impact from sporadic attacks. The training process of clients does not differ from the general training process, clients first receive the model $W_{r-1}$ from the server, then train the model $W_{r-1}$ with the local privacy training data for $E$ times and upload the trained model to the server. In other words, EC-LDA trades one round of training resources for one round of label distributions of client-side private training data.
IV-3 Extend to Link Prediction